# The StudyChat Dataset: Student Dialogues With ChatGPT in an Artificial Intelligence Course

### v2

@@ 4 Results and Analysis
In this section, we report on the overall trends we observe in the StudyChat dataset. We also analyze how students’ interactions with ChatGPT relate to course performance via a regression analysis.
Table 2: Distribution of dialogue act labels in StudyChat.
Conceptual Questions Contextual Questions Writing
Programming Language 219 Assignment Clarification 561 Write Code 617
Python Library 373 Code Explanation 783 Write English 163
Computer Science 113 Interpret Output 74 Conversion 207
Programming Tools 98 Other 1882 Summarize 64
Other Concept 352 Other 380
Total 1155 Total 3300 Total 1431
Verification Context Editing
Verify Code 237 Assignment Information 197 Edit Code 183
Verify Report 49 Error Message 253 Edit English 103
Verify Output 84 Code 268 Other 0
Other 44 Other 311
Total 414 Total 1029 Total 286
4.1 Dialogue Act Analysis
Table 2 shows the distribution of dialog acts in StudyChat. Off-Topic (173 counts) and Misc (68 counts) categories are relatively infrequent and not included in the table. The most common categories of dialogue acts are Contextual Questions, specifically the Assignment Clarification and Code Explanation acts (excluding than the generic dialogue act of Other). The relatively high frequency of contextual questions suggests that students are usually using the LLM to engage with and better understand the assignments, instead of simply copy-pasting LLM solutions. These observations are somewhat similar to those in Ma et al. (2024): students mostly tend to ask questions to the LLM rather than using it to directly complete assignments on their behalf. However, in our study, students were not discouraged from using the LLM to help co-author their code which may explain why in our study Writing > Write Code is still a very frequently occurring dialogue act.
4.2 Course Outcome Regression Analysis
We now investigate the relationship between student interactions with ChatGPT and their course performance via a multiple linear regression analysis. Our primary goal is to compare how different types of student-LLM interactions correlate with scores on items where ChatGPT was allowed and items where it was prohibited.
4.2.1 Feature selection
We select independent variables from two sources, the dialogue act labels and a pre-assessment survey given to students at the beginning of the course. We normalize all 42 total features against the maximum value across students, resulting in feature values in the range $[0,1]$. The dialogue act features are simply the counts of each type of dialogue act across each student’s utterances, and the student’s total number of utterances. In the survey, we ask students a series of questions about their background and prior experience, such as whether or not they had an prior experience with AI or machine learning. Then, on a 5-point Likert scale, we ask students to report: their confidence in three pre-requisite areas, including Python, Discrete Math, and Probability; how often they use LLM tools (from “Never” to “Everyday”); and how important they think LLM tools help their success in computer science.
Table 3: Regression coefficients for features with p-value $<0.2$.
Category Feature $\mathbf{\beta}$ p-value
ChatGPT Allowed
Provide Context Assignment Information 0.2249 0.03903
Conceptual Questions Other Concept 0.1111 0.04452
Contextual Questions Python Library 0.1178 0.15625
Verification Other -0.1366 0.18216
ChatGPT Prohibited
Provide Context Error Message -0.2937 0.02884
Provide Context Other 0.1904 0.05377
Pre-Survey Question Discrete Math Confidence 0.1557 0.06331
Conceptual Questions Programming Tools -0.1441 0.06613
Writing Request Code/Data Conversion 0.1247 0.06963
Contextual Questions Programming Language 0.1377 0.08329
Editing Request Edit Code 0.3072 0.18720
4.2.2 Regression Results
We fit two separate multiple linear regression models: The first model predicts student performance on the assignments and final group project, where using ChatGPT is allowed. The second model predicts student performance on quizzes and exams, where ChatGPT is prohibited. The dependent in both models is the aggregated score, normalized across all students in the dataset, also in a range $[0,1]$. We acknowledge that the StudyChat dataset only includes the scores of a subset of students who consented to share their assignment submissions, which somewhat limits the reproducibly of these results.
We observe an $R^{2}$ of 0.4136 on scores where ChatGPT usage is allowed and an $R^{2}$ of 0.5031 on scores where it is prohibited. We detail the most statistically significant features for each regression in Table 3. For items where ChatGPT usage is allowed, two features are statistically significant (p-value $<0.05$): Context > Assignment Information and Conceptual Question > Other. Both are positively correlated with student performance. We review 50 samples of utterances with these labels to explore trends in student behavior. For the Context > Assignment Information label, we find that most of these samples are instances of the student directly inserting all or parts of the instructions from the programming assignment into the chatbot. Therefore, it is unsurprising that this behavior would correlate with higher scores on the corresponding assignment, since it indicates the students are directly seeing, and perhaps copying ChatGPT’s solutions to the assignment. On the other hand, when reviewing the Conceptual Question > Other samples, we observe interactions where students ask questions that are tangential to the course, on topics such as statistics, ethical computing, computer networking, C++ programming libraries, and more. One explanation for the correlation between this feature and student performance is that it is a proxy for student interest and general curiosity about computer science, which may translate to better course engagement and completion.
In contrast, for quizzes and exams where ChatGPT usage is prohibited, only one feature, Context > Error Message, is statistically significant and it negatively correlates with performance. This observation suggests that students who tend to use ChatGPT to debug and fix error messages perform worse on tests. This finding seems counterintuitive: the course examinations do not ask any Python or library related questions and are focused on AI concepts and algorithms, so a student’s use of an LLM to help debug compiler errors seems unrelated. One possible explanation is this observation indicates student over-reliance on ChatGPT in completing the programming assignments, which may result in a lack of engagement with the core algorithms and concepts present in the assignment.
Extending our observations to borderline significant features (p-value $<0.07$), we see that Conceptual Questions > Programming Tools is also negatively correlated with exam performance, which could further support the hypothesis above of student over-reliance on ChatGPT. On the other hand, there are a number of borderline features that correlate with positive performance, such as Context > Other. When investigating these utterances, we see that these utterances show active student engagement in co-programming with the LLM. In these cases, unlike in the Context > Assignment Information feature, which is direct copy-pasting, the student is actively describing the requirements of the assignment in their own words to provide further information to the LLM. This behavior indicates the student’s active understanding of the assignment and underlying concepts, as they articulate what is needed to make progress on the assignment with a conceptual understanding of what they are doing. To put these features in perspective, students’ self-reported confidence in discrete math has a clear connection to exam performance, since data structure and algorithm knowledge are important background to the concepts covered in exams. However, this confidence feature is just as explanatory as the dialogue features and similarly significant. Finally, the Writing > Conversion feature is difficult to explain. In one of the assignments, students are directly asked to convert a table from one data format to another with ChatGPT, in order to explore the text generation capabilities of LLMs. In our review of the data, we only see cases of students completing this part of the assignment and no other conversions, which only indicates that they are following the assignment instructions.
Our initial regression analysis of StudyChat provides some insights, albeit noisy ones, into the relationship between student behavior with LLMs and course performance. We see that students copy-pasting behaviors correlate with better assignment grades, but we do not observe a negative correlation between this behavior and exam performance, but instead see this correlation with help-seeking behaviors (tool assistance and debugging). These observations suggest that there could be student behavior patterns with LLMs that could potentially be monitored and detected in real time, enabling teachers to intervene and help students who are exhibiting patterns that correlate with at-risk performance.
5 Future Research Directions
In this work, we provide an initial analysis of the StudyChat dataset and surface-level insights into this data. In this section, we outline additional avenues for possible future work on this data and the kind of research questions that can be asked and possibly answered.
First, since this dataset includes time stamps on each student interaction with the LLM, we can track student behavior over time, both within an assignment and across assignments over the course of the semester. For example, we can study the variation in student engagement levels over each assignment: whether student activity drops off over the semester, and if there are any correlations between disengagement and course performance. In addition, we can study whether increased engagement with an LLM leads to better course performance and assignment solutions.
Second, since this dataset also includes student assignment submissions, we can further compare these submissions against student utterances when talking to the LLM. For example, we can study how closely final student submissions mirror the exact LLM output, to quantify whether students are simply copy-pasting or using LLM output for helpful suggestions. We can also investigate how such behavior correlates with course performance. This analysis may inform future content developers to develop homework questions that cannot be answered by solely relying on LLM output or perhaps integrate prompt strategies into the assignments that teach effective usage of LLMs and common pitfalls.
Third, since our initial labeling schema has clear room for improvement, we can investigate more accurate and fine-grained methods of classifying student utterances. For example, we can use span-level tagging to isolate parts of student utterance that are of particular interest, which may enable a more thorough discourse analysis and findings. Moreover, one may investigate n-gram dialogue acts and even study how student intent evolves over a conversation, which may reveal more insights into how students react to different LLM output. These insights can potentially inform future work on aligning LLMs with pedagogical principles.
Finally, our initial findings of dialogue acts which correlate with negative student performance is only a beginning into a proper investigation into the potential negative learning side-effects of LLM usage. LLMs are not oracles, and when they hallucinate information they could confuse students or worse, potentially instill misconceptions. A more thorough investigation into the accuracy of the statements made by the chatbot is certainly warranted, and if false statements made by the LLM lead to more buggy code or worse course outcomes.
6 Conclusions
@@
In this work we present the StudyChat dataset, a collection of over 1,000 conversations between students and ChatGPT in an upper-division, university-level artificial intelligence course. We systematically create a dialogue act labeling schema which we applied to all utterances in the dataset via LLM labeling and validate this labeling with human agreement. We analyze the resulting labeled dataset via a regression analysis of student behavior, which suggests that students tend to use the model largely for question answering more than any other activity and, in some cases, this correlates to better course performance, but other help-seeking behaviors correlate with lower student performance.

### v3

@@ 4 Results and Analysis
In this section, we study the relationship between student interactions with the LLM chatbot and their course outcomes.
Figure 2: Average assignment scores (LLM-supported) vs. exam scores (No-LLM) for both semesters, grouped by interaction level: bottom 10% (red), middle 80% (blue), and top 10% (green). Grey dashed line shows equal outcome and red line shows data trend; top users cluster at higher scores and bottom users show wider spread and more low outliers.
4.1 Dialogue Acts as Assessment Predictors
We postulate that student interactions with LLMs during assignments will have an observable influence on student course outcome. To study this hypothesis, we run a series of regression models and measure the varying effects of adding LLM usage features of increasing specificity. First, we use a single feature of the total count of utterances, then a series of features 8 corresponding to the broad DA label counts, and finally a series of 31 features corresponding to the specific DA label counts. For all models, we use a dependent variable of the score of an assessment normalized to the range $[0,1]$. In this analysis, we define assessment as one of the 7 assignments (a1-a7) or 3 course exams (e1-e3). For our baseline model, we use one independent variable, the running average of prior assessment scores in the range $[0,1]$, a well studied predictor of post assessment score hallberg2018pretest. We run each model 18 times, for 9 assessments across both semesters, excluding a1 which has no prior assessment. For the baseline model, the mean $R^{2}$ values for all regressions are $0.17343$ and $0.17281$, for the fall and spring semesters, respectively.
For our first model, we add one additional independent variable to the linear regression model: the total count of all utterances inclusive of the current assignment. For exams, we sum the utterances for all assignments that were due before the exam. For these regressions, the mean $R^{2}$ was $0.1847$ for the fall and $0.1860$ for the spring, with mean $R^{2}$ improvements of $0.0064$ and $0.0178$, respectively. We run F-tests for each improvement and we find that for one assessment, spring semester a5, the addition of the usage count independent variable leads to a statically significant improvement in explained variance, with an increase of $0.0242$, $F=5.82$, $p=0.02$. In the other cases, we found no statistically significant improvement.
For our next two models, we explore the influence of incorporating DA information into the regression. We expand our independent variables to include features based on the two levels of DA behavior labels, using 8 broad-labeled DAs or 31 specific-labeled DAs respectively. Specifically, we augment the baseline regression by including the counts of utterances for each of the DA labels present in the assessment target. In the same manner as above, for exams, we aggregate all assignment-specific DA counts that were due prior to the exam. With the inclusion of broad DA label counts, we observe that this inclusion has a mean $R^{2}$ of $0.2529$ for the fall and $0.2288$ for the spring semester, with a mean improvement of $0.0746$ and $0.0606$, respectively. These $R^{2}$ values indicate that DA features contribute explanatory power but F-Tests conducted for each regression do not indicate results that are statistically significant, with the highest $R^{2}$ improvement of $0.19$ and resulting F-Score of $F=1.81,p=0.09$. Repeating the same experiment with the specific DA label count features, we observe this inclusion increases the mean $R^{2}$ to $0.4868$ for the fall semester and $0.4012$ for the spring semester, with a mean $R^{2}$ increase of $0.3086$ and $0.2330$ respectively. Similar to the broad features, these $R^{2}$ values indicate that DA features contribute explanatory power but F-Tests conducted for each regression do not indicate results that are statistically significant. One assessment, a4 from the fall semester, has a borderline significant improvement of $0.4$ with an F-Score of $F=1.77,p=0.05$.
Our findings in these regression analyses have mixed results in support of our hypothesis. The analysis indicates that the explanatory power of a linear regression model increases significantly when moving from a simple count of utterances to a more nuanced, qualitative representation based on DAs. However, the relative explanatory power compared to a simple, prior performance-based explainer of student outcome is only statistically significant when the utterances are represented as aggregate counts, which may simply be a confounder for student engagement or participation. This discrepancy suggests that the DA-based models are likely overfitting the data or are statistically underpowered by limited sample size. While the relationship between dialogue patterns and outcome likely exists, a more sophisticated feature selection method and modeling approach would be required to build a model that is highly predictive and robust in explaining student assessment course outcomes.
We further analyze the more granular relationships between DA and assessment outcomes by examining significant coefficients in our regression models. Specifically, for the broad and specific DA regressions models, we run F-Tests to compute the significance of the coefficients for each target. Across the 36 regressions, 20 coefficients were found to be statistically significant, which we report in Table 4. For the fall semester, we observe significant coefficients only from specific-level DA regressions and only for assignment targets. Most coefficients are negatively correlated with assignment scores and largely fall into other specific DA labels, with the only positive significant correlation attributed to an off-topic DA. One possible explanation for this observation is, compared to other DAs, these labels occur relatively infrequently in the distribution, and the linear regression was likely skewed by a small subset of student behavior patterns which were uncommon.
The spring semester significant coefficients are mostly distinct from the Fall, and seem to have potentially more meaningful pedagogical takeaways. One consistent theme across both semesters is that Contextual Questions acts are negatively correlated with assignment outcomes. Unique to the Spring regressions are statistically significant coefficients when the target is an exam score and the inclusion of broad-level DA features. We observe the broad feature for Conceptual Questions and Editing Request as positively correlated with e1 outcomes, indicating a learning-focus usage of LLMs being positively correlated with assessment outcome. Conversely, the Contextual Questions broad level DA was negatively correlated with Assignment 5 outcome. Since the Contextual Questions label corresponds to student questions are related to the specifics of the current conversation, this negative correlation is perhaps an indicator of student confusion. On the other hand, the Conceptual Questions label indicates a student asking a general knowledge question to the chatbot instead of something specific to the course. One possible takeaway from this finding is that educators can foster student use of LLMs for general knowledge and well-established concepts, but discourage asking for course-specific help. These suggestions may help to strike a balance for educators in fostering a modern technology-driven classroom, while preventing students from developing misconceptions from inaccurate chatbot responses.
In investigating the specific DAs in the spring semester, we see Assignment 5 as a consistent target and also observe specific contextual question features as having negative correlation, with the exception of Contextual Question - Code Explanation, which has a slight positive correlation. For the spring dataset regressions, we also see different specific types of Conceptual Questions being mostly positively correlated with scores. One notable exception is that Conceptual Question - Mathematics is a strong negative indicator of a6 score. In this assignment, students were tasked with building an n-gram language model, which involved computing probability tables and deriving equations for the model. This correlation suggests that students who relied heavily on the model to understand these equations and derivations did not score well, potentially deriving incorrect solutions from the model or receiving inaccurate guidance. This observation suggests that LLM conceptual knowledge and explanations are inconsistent and that LLMs, or at least gpt4o-mini, are poor at explaining or deriving math equations. This finding suggests that educators should advise students to be cautious with LLM explanations and to inform them that while LLMs may have accurate general knowledge that the procedural derivations of these models should be carefully interpreted and validated.
4.2 Low and High Usage Users
To understand how the level of user interaction influences outcome, we examine students in the bottom 10 %, middle 80% and top 10 % in terms of total utterances over the course of the semester. We will refer to these groups as low, medium and high interaction users. In the Fall dataset, the low interaction users averaged 8.6 utterances per student, and the high interaction users averaged 231.8 utterances per student. In the Spring, the low interaction averaged only 3.2 utterances per student, and the high interaction users averaged 284 utterances per student.
To assess how interaction levels relate to student outcomes, we run multiple linear regression models. Identically to the previous section, we add control independent variable to each model (the running average of prior assessment scores), use a dependent variable that is the normalized score of the assessment in the range $[0,1]$, and run this configuration for our baseline regressions. For group membership, we include three additional independent binary variables which indicate whether the student was in the low, medium, or high interaction group and conduct these regressions separately for the Fall and Spring semesters.
The baseline mean $R^{2}$ scores across all regressions are 0.1783 for the fall semester and 0.1682 for the spring semester. Adding the additional group features increases the mean $R^{2}$ to 0.2032 and 0.1858, respectively. This additional categorical variable yields a borderline significant improvement in explained variance for one assessment, Spring e1, $F=4.08$, $p=0.05$. On this assessment, the average scores for the low interaction, medium interaction, and high interaction users are 77.8%, 81.0%, and 88.1%, respectively. Corroborating with our finding in the previous section it appears that students with high usage, focused on conceptual questions on early assignments tend to perform better on the first exam. While the grouping-level features are statistically noteworthy, the modest gains suggest that usage patterns have limited explanatory power for overall course outcome.
We compare each group by plotting individual students’ average assignment scores and average exam scores, as shown in Figure 2. We observe the minimum average assignment score (92.53), and the minimum average exam scores (72.33) for the high interaction users across both semesters are higher than the minimum of the other two groups. Moreover, the variance in the average exam and assignment scores is smaller for high interaction users. For the remaining 90% of users (low and medium interaction users), there was no substantial difference in minimum scores or variance between groups. Our analysis demonstrates that although interaction level is not a dominant predictor of outcome, high engagement with the LLM is associated with reduced variability in course outcomes and higher minimum course outcome. This pattern suggests that encouraging consistent and meaningful LLM usage, particularly among students who need additional tutoring and guidance, may help stabilize performance and raise lower-end outcomes, even if the overall effect on average scores is modest.
4.3 Clustering Students by Behavior
Figure 3: PCA of student DA features for Fall (left) and Spring (right). Points show students colored by K-means cluster. PC1 separates technical question from writing behaviors; PC2 separates context-seeking from code-oriented behaviors.
To investigate whether distinct patterns of student behavior emerge, we apply unsupervised clustering to behavioral interaction features. We follow an approach similar to Brender et al. (2024), forming feature vectors that normalize each student’s DA utterance types by their total count of utterances. In total, there are 39 features, with the 8 broad DA features and 31 specific features. These vectors represent each student’s distribution of DA labels over the course of the semester. We use K-means clustering on these feature vectors to group students based on their behavioral profiles, applying the method separately to Fall and Spring semester data. The optimal number of clusters, $k=4$, is determined using the elbow method and silhouette analysis.
We visualize these groupings with Principal Component Analysis (PCA) in Figure 3, reducing the behavioral feature space to two components that captured the largest sources of variation (Fall: PC1 = 12.6%, PC2 = 9.6%; Spring: PC1 = 11.1%, PC2 = 9.3%). PC1 contrasts technical question–oriented usage (conceptual questions, Python library, programming language queries) with writing-focused (writing code) and context-providing behaviors; PC2 contrasts off-topic behavior (chit chat, greeting) with writing (write english, write code) and conceptual questions. Overlaid cluster boundaries and labels in the plots highlight these behavior styles and their degree of separation.
We report the distribution of DA for each cluster across for the fall and spring in Table 5 and  6 respectively. While the exact makeup of each cluster varies across semesters, some general patterns emerge. For example, certain broad labels, such as writing request and contextual question labels, are prevalent across multiple clusters, while others (e.g., off-topic or conceptual questions) are concentrated in specific clusters. Notably, some clusters represent high engagement (e.g., students with over 100 total utterances), while others capture minimal or off-topic usage. These differences suggest a spectrum of interaction styles, from consistent academic engagement to sparse or misaligned LLM usage. We see two recurring patterns of cluster assignment: groups with frequent writing requests and groups with frequent question-asking requests.
We examine the differences in the similar clusters in the fall semester and find patterns emerge. For the writing-request clusters, students in cluster 0 tend to ask the chatbot to write code while students in cluster 3 display a mixture of code writing and english writing requests. This observation suggests that students in cluster 3 lean on the LLM not only for help in implementing the requested program, but also in synthesizing the results and writing assignment reports. For the question-asking clusters, students in cluster 1 tend to ask conceptual questions, mostly on and Python-related questions and computer science topics, while students in cluster 2 ask a mixture of conceptual questions and contextual questions of a much wider spread of topics. This observation suggests that students in cluster 1 use the LLM for understanding the tools needed to complete the assignments, while students in the later category use the LLM as a more general purpose tutoring system. In examining the average assignment and exam scores between these clusters, we see that all clusters have comparable average assignment scores. However coding-question groups trends towards a higher average score on the exams $89\pm 7.3$ while report-writers tend to have a marginally lower average score $83.8\pm 11.7$. These observations suggest that students who ask targeted questions of the LLM for specific programming and coding concepts while working on assignments tend to enter exams with a better understanding of the material than students who request extensive writing assistance, which may circumvent the learning objectives of the assignments.
To further explore these trends in assessment scores, we again perform a regression analysis. We use the same baseline model, a single independent variable, running average of prior assessment scores, to dependent variable of assessment grades. Similar to the usage interaction analysis, for cluster membership we include four additional binary variables, each associated with a group to assess its additional explanatory power. We find that adding the cluster variable improved the predictive performance marginally. For the Fall semester, the mean $R^{2}$ increased from 0.1783 to 0.2152; for Spring, it improved from 0.1682 to 0.1844. We conduct F-tests for each regression and find that no indications of statistically significant improvement in explanatory power, with the highest Resulting F-score of F=2.67, p=0.06. These results demonstrate that while this behavioral clustering does provide an improvement, similar to the usage grouping, we cannot definitively conclude this clustering strategy is statistically meaningful in terms of understanding course outcomes. However, while the clusters are not definitive predictors, they do offer interpretable groupings of student behavior that correlate with distinct behavior styles ranging from frequent, academically aligned interactions to sparse or off-topic usage. This diversity highlights that engagement is multidimensional, varying not only in frequency, but also type of behavior. These clusters provide quantitative insights into patterns of student LLM usage which may prove helpful in educating students to use these tools in a learning-positive way.
5 Conclusions, Limitations, Future Work
@@
In this work, we present the StudyChat dataset, a collection of over 16,000 interactions between students and ChatGPT in a university-level artificial intelligence course. We systematically create and apply a dialogue act labeling schema applied to this dataset via LLM labeling, validate this labeling with human agreement, and conduct a regression and clustering analysis to analyze trends of student behavior as they relate to course outcomes.
While our analysis offers new insights into how students interact with LLMs, the takeaways from our work have limitations. First, our dataset was drawn from a single upper-division AI course, which limits how well the findings generalize. Second, participants in this study were aware that their conversations were recorded and thus their behaviors may have been altered and subject to the Hawthorne Effect landsberger1958hawthorne.
Our findings from this exploration and the resulting dataset open many avenues for future work. First, since the StudyChats dataset also includes student assignment submissions, we can further compare these submissions against student utterances when conversing with the LLM which may reveal further behavioral insights and trends. Second, our initial labeling schema and labeling methodology could be refined further, to better cover overlapping labels and missing dialogue acts with more accurate and fine-grained methods of classifying student utterances. Finally, these insights can be taken back into a classroom setting where we could create LLM-based tutoring tools that detect and discourage behaviors associated with negative course outcomes.
@@ Acknowledgments
Appendix A Appendix
Table 3: Distribution of dialogue act labels in StudyChat.
Conceptual Questions Contextual Questions Writing
Python Library 1788 Assignment Clarification 1076 Write Code 2534
Programming Language 1271 Code Explanation 950 Write English 1340
Computer Science 1178 Interpret Output 513 Conversion 347
Programming Tools 484 Other 399 Summarize 52
Mathematics 155 Other 63
Other Concept 323
Total 5199 Total 2938 Total 4336
Verification Context Editing
Verify Code 420 Assignment Information 434 Edit Code 213
Verify Report 100 Error Message 979 Edit English 197
Verify Output 77 Code 696 Other 8
Other 50 Other 919
Total 647 Total 3028 Total 418
Misc Off Topic
Other 30 Chit-Chat 102
Total 30 Gratitude 35
Greeting 84
Other 29
Total 250
Table 4: Summary of statistically significant regression coefficients ($p<0.05$) for predicting assessment outcomes from interaction features, grouped by semester and feature set. Only significant results are listed.
Semester Feature Set Feature Name Target Coefficient $p$ $n$
F24 Broad No significant coefficients
Specific Contextual Questions: Code Explanation Assignment 3 0.0065-0.0065 $0.021$ $72$
Contextual Questions: Other Assignment 3 0.0156-0.0156 $0.034$ $72$
Off Topic: Chit-Chat Assignment 3 $0.0235$ $0.043$ $72$
Provide Context: Code Assignment 4 0.0426-0.0426 $0.001$ $66$
Verification: Other Assignment 4 0.0291-0.0291 $0.007$ $66$
Conceptual Questions: Other Concept Assignment 5 0.0575-0.0575 $0.025$ $64$
Writing Request: Other Assignment 5 0.0590-0.0590 $0.033$ $64$
S25 Broad Conceptual Questions Exam 1 $0.0133$ $0.024$ $104$
Editing Request Exam 1 $0.0296$ $0.017$ $104$
Contextual Questions Assignment 5 0.0020-0.0020 $0.005$ $72$
Specific Conceptual Questions: Python Library Exam 1 $0.0234$ $0.010$ $104$
Contextual Questions: Assignment Clarification Assignment 3 0.0034-0.0034 $0.021$ $103$
Conceptual Questions: Other Concept Assignment 3 $0.0095$ $0.029$ $103$
Writing Request: Conversion Exam 2 $0.0028$ $0.045$ $119$
Contextual Questions: Assignment Clarification Assignment 5 0.0072-0.0072 $0.039$ $72$
Contextual Questions: Code Explanation Assignment 5 $0.0009$ $0.044$ $72$
Contextual Questions: Other Assignment 5 0.0075-0.0075 $0.017$ $72$
Conceptual Questions: Other Concept Assignment 5 $0.0016$ $0.019$ $72$
Conceptual Questions: Mathematics Assignment 6 0.1062-0.1062 $0.002$ $74$
Off Topic: Gratitude Exam 3 0.0337-0.0337 $0.025$ $119$
Figure 4: Usage over time organized by assignments for the Fall 2024 semester
Figure 5: Usage over time organized by assignments for the Spring 2025 semester
Figure 6: Box plot showing the distribution of total utterances per student for each conversation topic where outliers are shown with cross markers.
Table 5: Top 10 Dialogue Acts per Cluster – F24 Semester
Cluster Users Dialogue Act % Occurrence
Cluster 0 25 Broad Writing Request 36.66
Broad Provide Context 24.73
Specific Writing Request Write Code 24.44
Broad Conceptual Questions 17.43
Broad Contextual Questions 14.98
Specific Writing Request Write English 8.63
Specific Provide Context Error Message 8.05
Specific Provide Context Other 7.37
Specific Contextual Questions Assignment Clarification 5.94
Specific Provide Context Code 4.82
Cluster 1 16 Broad Conceptual Questions 63.02
Specific Conceptual Questions Python Library 26.31
Specific Conceptual Questions Programming Language 15.64
Broad Writing Request 15.32
Specific Conceptual Questions Computer Science 13.09
Specific Writing Request Write Code 11.32
Broad Contextual Questions 10.84
Broad Provide Context 8.24
Specific Conceptual Questions Programming Tools 4.04
Specific Contextual Questions Code Explanation 3.98
Cluster 2 27 Broad Conceptual Questions 36.20
Broad Contextual Questions 25.88
Broad Writing Request 17.89
Specific Conceptual Questions Python Library 15.97
Broad Provide Context 9.85
Specific Contextual Questions Code Explanation 9.39
Specific Writing Request Write Code 9.35
Specific Contextual Questions Assignment Clarification 8.85
Specific Conceptual Questions Computer Science 7.67
Specific Conceptual Questions Programming Language 6.46
Cluster 3 16 Broad Writing Request 50.25
Specific Writing Request Write Code 17.75
Broad Conceptual Questions 17.18
Specific Writing Request Write English 16.79
Specific Writing Request Code/Data Conversion 14.47
Broad Provide Context 10.72
Broad Contextual Questions 8.14
Specific Conceptual Questions Computer Science 8.13
Broad Verification 6.61
Specific Verification Verify Code 5.42
Table 6: Top 10 Dialogue Acts per Cluster – S25 Semester
Cluster Users Dialogue Act % Occurrence
Cluster 0 5 Broad Provide Context 46.85
Specific Provide Context Other 23.95
Broad Writing Request 14.63
Broad Conceptual Questions 13.12
Broad Off Topic 10.88
Specific Off Topic Greeting 10.63
Specific Provide Context Code 10.11
Broad Contextual Questions 9.95
Specific Provide Context Error Message 8.49
Specific Writing Request Write Code 7.95
Cluster 1 68 Broad Writing Request 44.63
Specific Writing Request Write Code 22.28
Broad Contextual Questions 18.46
Broad Provide Context 17.04
Specific Writing Request Write English 12.99
Broad Conceptual Questions 10.75
Specific Contextual Questions Assignment Clarification 10.30
Specific Writing Request Code/Data Conversion 7.13
Specific Provide Context Error Message 5.00
Specific Provide Context Other 4.76
Cluster 2 43 Broad Conceptual Questions 43.55
Broad Writing Request 21.25
Broad Contextual Questions 17.97
Specific Conceptual Questions Python Library 15.67
Specific Writing Request Write Code 12.11
Specific Conceptual Questions Computer Science 11.29
Broad Provide Context 10.44
Specific Conceptual Questions Programming Language 10.19
Specific Contextual Questions Assignment Clarification 6.31
Specific Contextual Questions Code Explanation 6.27
Cluster 3 3 Broad Off Topic 41.82
Specific Off Topic Chit-Chat 29.09
Broad Contextual Questions 26.06
Broad Conceptual Questions 23.03
Specific Conceptual Questions Other Concept 20.00
Specific Contextual Questions Other 20.00
Specific Off Topic Greeting 12.73
Broad Provide Context 3.03
Broad Verification 3.03
Broad Writing Request 3.03